OpenStudy (anonymous):

Graph the function. Then use your graph to find the indicated limit. lim f(x) = 5x - 3, x --> 5, f(x) A. 5 B. 25 C. 2 D. 22

OpenStudy (anonymous):

Looks like D to me as x approaches 5 we may as well round it to 5 5(5)-3 = 25-3 = 22

OpenStudy (anonymous):

thank you